3. Solve the following problem from t 0 to 1 with h-1 using 3rd order RK method: dx dt dy dt bay where (0)-4 and x(0)- 0.
Answer #1

da lt dt ブ丨こプ叶 2-,% 0.5ノ6,34i f ( , o+ 2 @as)-9, aws.ast。) 94 -20 692861 - o+ 25846) = n+8206930799 ズ1-7(リ= 3 82799, 6667